OpenStudy (anonymous):

Solve g(x) if g'(x) = sin 3x + cos 2x and g(pi) = 2 I get the following: G(pi) = (cos 3pi / 3) - (sin 2pi / 2) + C = 2 G(pi) = (cos pi) - (sin pi) + C = 2 G(pi) = -1 - 0 + C = 2 C = 3 In the correct answer, C = 5/3

OpenStudy (mertsj):

\[g(x)=\frac{1}{2}\sin 2x-\frac{1}{3}\cos 3x+C\]

OpenStudy (mertsj):

\[\frac{1}{2}\sin 2\pi-\frac{1}{3}\cos 3\pi+C=2\]

OpenStudy (mertsj):

\[0-\frac{1}{3}(-1)+C=2\]

OpenStudy (mertsj):

\[C=2-\frac{1}{3}=\frac{5}{3}\]

OpenStudy (anonymous):

Thank you!

OpenStudy (mertsj):

yw